OpenStudy (anonymous):

please help me i only have two questions left to finish Nate and Haley ride to school 1.6 miles north and then 3 miles west. If a straight line is drawn from their house to the school, how far is the distance? Round your answer to one decimal place. 2nd question If the midpoint of two points is (22,18) and one point is (10,30), find the other point.

OpenStudy (anonymous):

1.6 and 3 are the sides of a right triangle and the direct route is the hypotenuse so use Pythagorus. distance = sqrt(3^2+1.6^2) = 3.4 miles Second one (x+10)/2 =¨22 so x is 34 and (y +30)/2 = 18 so y is 6 so (34,6)

OpenStudy (anonymous):

thank you so much